# Preserving the Hermiticity of the One-Body Density Matrix for a Non-Interacting Fermi Gas

### v2

@@ Abstract
The one-body density matrix (ODM) for a $d$-dimensional zero temperature non-interacting Fermi gas can be approximately obtained in the semiclassical regime through different $\hbar$-expansion techniques. One would expect that each method of approximating the ODM should yield equivalent density matrices which are both Hermitian and idempotent to any order in $\hbar$. However, the Kirzhnits and Wigner-Kirkwood methods do not yield these properties, while the method of Grammaticos and Voros does. Here we show explicitly, for arbitrary $d\geq 1$-dimensions through an appropriate change into symmetric coordinates, that each method is indeed identical, Hermitian, and idempotent. This change of variables resolves the inconsistencies between the various methods, showing that the non-Hermitian and non-idempotent behaviour of the Kirzhnits and Wigner-Kirkwood methods is an artifact of performing a non-symmetric truncation to the semiclassical $\hbar$-expansions. Our work also provides the first explicit derivation of the $d\geq 1$-dimensional Grammaticos and Voros ODM, originally proposed by Redjati et al. [1] based on their $d=1,2,3,4$ expressions.
Keywords: Density-functional theory, semiclassical, $\hbar$-expansion, Hermitian, ultracold Fermi gas
@@ 1 Introduction
Density functional theory (DFT) is a powerful tool used to solve quantum many-body problems. It has a wide range of applications in different branches of science, see [2, 3, 4, 5, 6, 7] and references therein for some examples. Particularly, Hohenberg-Kohn-Sham DFT allows one to map $N$ interacting fermions to $N$ non-interacting fermions under the influence of an effective one-body potential $v_{\mathrm{eff}}(\mathbf{r})$ [8]. In $d\geq 1$-spatial dimensions the total-energy of the system can then be expressed as a unique functional of the one-body density matrix (ODM), $\rho$, by
$E[\rho]=T_{0}[\rho]+\int d^{d}rv_{\mathrm{eff}}(\mathbf{r})\rho(\mathbf{r}).$ (1.1)
$T_{0}$ is the kinetic energy functional, and $v_{\mathrm{eff}}[\rho]$ can be split into 1) a combined classical plus quantum interaction term $E_{\mathrm{int}}[\rho]$, and 2) a term accounting for the presence of any external potential $v_{\mathrm{ext}}(\mathbf{r})$:
$v_{\mathrm{eff}}(\mathbf{r})=\frac{\delta E_{\mathrm{int}}[\rho]}{\delta\rho}+v_{\mathrm{eff}}(\mathbf{r}).$ (1.2)
@@
In practice Hohenberg-Kohn-Sham theory hinges on finding approximate expressions for the above density functionals, where perhaps the simplest is provided by the local-density approximation which assumes a simple form of $\rho$ that is locally valid even for inhomogeneous systems [9, 10]. Although this approximation is an excellent theoretical starting point it often fails to capture the complete physics of a given non-interacting Fermi gas since it neglects the effects of inhomogeneities. For non-uniform systems the exact ODM (and therefore exact density functionals) are in general not known, but density gradient corrections to the local-density approximation can be computed through various techniques [8, 10, 11, 12, 13]. A subset of these techniques include different semiclassical $\hbar$-expansions, three popular methods being the Kirzhnitz method (KODM) [14, 15, 16, 17], the Wigner-Kirkwood method (WKODM) [18], and the method of Grammaticos and Voros (GVODM) [19]. In reference [20] the KODM was derived up to second order in $\hbar$ for arbitrary $d\geq 1$-dimensions. Furthermore, in reference [10] the $d\geq 1$-dimensional WKODM method is discussed although the ODM is not explicitly derived, and in references [1, 21, 22] the GVODM is derived to second order in $\hbar$ for $d=1,2,3,4$-dimensions. Although in each case the ODMs were all derived for a zero temperature non-interacting Fermi gas in the presence of a position-dependent external potential, they can still be used to account for the effects of finite temperature and internal interactions, see references [22, 23].
@@
The KODM and WKODM methods are Hermitian $*\rho(\mathbf{r},\mathbf{r}\prime)=\rho(\mathbf{r}\prime,\mathbf{r})^{*}$ and idempotent $\hat{\rho}^{2}=\hat{\rho}$ if all terms in the $\hbar$ expansion are accounted for, but they are not guaranteed to be if their expansions are truncated at a given order in $\hbar$. In particular, within reference [20] the $d\geq 1$-dimensional KODM is truncated to second order in $\hbar$, and is seen to be manifestly non-Hermitian and non-idempotent. In contrast, the GVODM preserves Hermiticity and idempotency regardless of whichever order in $\hbar$ the semiclassical expansion is truncated at. These issues are commented on in references [1, 21, 22]. The various ODMs should all be equivalent since they must yield the same physical properties of the Fermi gas, e.g., the total energy. Therefore it is unsatisfying that the GVODM is manifestly Hermitian and idempotent while the KODM and WKODM are not. Additionally, a $d\geq 1$-dimensional expression for the GVODM has not yet been explicitly derived. Although reference [1] proposed such an expression based on their $d=1,2,3,4$ results, a rigorous derivation was not performed.
@@
The structure of this paper is as follows: In Section 2, we show that because the KODM is formulated in terms of inherently non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$ from the beginning of its derivation, and it does not display the same symmetry as the GVODM which is formulated in terms of center of mass $\mathbf{R}$ and relative $\mathbf{s}$ coordinates, shown in equation 1.3. These symmetric coordinates are typically better suited to help a system display any inherent symmetries it may possess, since the center of mass between two particles remains invariant under translations, rotations, and reflections of the entire system, while the distance between any two particles remains invariant under translations and rotations of the system with respect to $\mathbf{R}$.
@@
We explicitly show in $d\geq 1$-dimensions that, upon the proper symmetrization procedure (change of coordinates) from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$, the KODM is identical to the GVODM up to second order in $\hbar$. Our starting point relies on the ability to express each term of the KODM as a specific inverse Laplace transformation before the coordinate transformation is undertaken. We show in A that in terms of non-symmetric coordinates, the WKODM is identical to the KODM via these inverse Laplace transformations [10, 24], allowing us to make the necessary change of coordinates. In Section 3, we address the fact that the GVODM originates from the non-Hermitian and non-idempotent WKODM method in terms of non-symmetric coordinates, even though it results in a Hermitian and idempotent ODM in terms of symmetric coordinates. Similar to Section 2, we show in $d\geq 1$-dimensions that by applying a symmetric WKODM method in terms of $(\mathbf{R},\mathbf{s})$ coordinates from the very beginning, the resulting WKODM is identical to the GVODM to second order in $\hbar$ as expected.
The symmetrization procedure for both the KODM and WKODM demonstrates that regardless of whichever method is used to derive the ODM of a zero temperature non-interacting Fermi gas, the same Hermitian and idempotent expression is obtained as one would expect. Our results can be shown to hold when higher order terms in the ODMs are accounted for, but, for simplicity, we restricted our analysis to second order in $\hbar$. This therefore resolves the Hermiticity and idempotency inconsistency between the various methods, showing that problems only arise when non-symmetric truncations of semiclassical expansions are taken. Additionally this explicitly shows that the predicted $d\geq 1$-dimensional expression for the GVODM proposed in reference [1] is in fact correct, as we exactly obtain their result through our symmetrization procedure for both the KODM and WKODM methods. To our knowledge, these issues have been overlooked in the semiclassical expansion methods utilized in DFT.
2 Hermitian form of the KODM in $d\geq 1$-dimensions
The $d\geq 1$-dimensional KODM for a non-interacting Fermi gas is given in non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$ by equation 2.1 [20].
$\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})=A_{0}+A_{1}+A_{2};$ (2.1)
$A_{0}$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z),$ (2.2)
$A_{1}$ $=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)},$ (2.3)
$A_{2}$ $=\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{r}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{r}}k_{F}^{2}\big{)}^{2}$
@@ 2 Hermitian form of the KODM in $d\geq 1$-dimensions
The local Fermi wavenumber is defined by $k_{F}=k_{F}(\mathbf{r})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{r})\big{)}^{1/2}$ where, for convenience, we have suppressed the $\mathbf{r}$ coordinate dependence of both $k_{F}$ and the $A_{1,2,3}$ terms. $\mu$ is the chemical potential which is equal to the local Fermi energy for our case of a zero temperature Fermi gas, and $z=k_{F}s$ with $s=|s|=\sqrt{\mathbf{s}\cdot\mathbf{s}}$. $J_{\nu}(z)$ is a cylindrical Bessel function of order $\nu$ [25], and $g$ is a fermionic degeneracy factor where $g=1$ when the Fermi gas is spin polarized, and $g=2$ when it is not, since each fermion will then have two possible spin states.
@@
Equation 2.1 was obtained by simplifying equation 17 from reference [20] using the Bessel function identities described in B, namely equations 2.1 and 2.2. Furthermore, we have collected the various terms of the ODM based on their order of $\hbar$ (or equivalently, their order of $\nabla_{\mathbf{r}}$). $A_{0}$ is the classical zeroth order in $\hbar$ term of the semiclassical expansion, also known as the Thomas-Fermi term. It is the only term which does not vanish in the local density regime, where effects due to inhomogeneities are sufficiently small such that $s$ is negligible compared to the scale over which the density changes, i.e., $s\approx 0$, yielding $\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})\approx A_{0}$ [21, 22]. Going beyond the local density approximation and including quantum density gradient corrections yields $A_{1}$ and $A_{2}$, proportional to $\hbar$ and 2superscriptPlanck-constant-over-$\hbar^{2}$, respectively.
@@
Equation 2.1 is manifestly non-Hermitian and non-idempotent as discussed in references [1, 21, 22]. However, by changing to center of mass $\mathbf{R}$ and relative coordinates $\mathbf{s}$ via equation 1.3, we show that equation 2.1 can be put into a manifestly Hermitian and idempotent form identical to the GVODM. The transformation from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ variables is non-trivial as we will show. Consequently, to simplify the process we will individually transform each $A_{1,2,3}$ term in equation 2.1 and sum the results in the end via equation 2.1.
Although reference [20] derives equations 2.1, 2.2, 2.3 and 2.4 for $d\geq 1$-dimensions in terms of non-symmetric coordinates using the commutator method of Kirzhnitz [14], they can equivalently also be derived in these coordinates via the WKODM method [10, 18, 24]. We show this to be true in A so that equations 2.2, 2.3 and 2.4 may alternatively each be expressed by the following inverse Laplace transforms, respectively.
$A_{0}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.5)
$A_{1}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.6)
$A_{2}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.7)
where $\beta$ is a complex parameter which is a conjugate variable to the chemical potential $\mu$, not to be confused with the inverse of the temperature $(k_{B}T)^{-1}$ [23]. $C(\mathbf{r},\mathbf{r}^{\prime};\beta)$ is known as the Bloch density matrix, which has been expanded to second order in $\hbar$ so that $C(\mathbf{r},\mathbf{r}^{\prime};\beta)\approx C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$. Similar to our descriptions of the $A_{1,2,3}$ terms, $C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ is the zeroth order in $\hbar$ classical (Thomas-Fermi) term, $C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ is the first order in $\hbar$ term, and $C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ in the second order in $\hbar$ term. These are given by equations 2.8, 2.9 and 2.10 and the details their origin is addressed in more detail within A by equations 1.1, 1.2, 1.3, 1.4 and 1.5 [10].
$C_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (2.8)
@@
$C_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{p}\Big{)},$ (2.9)
@@
$C_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\bigg{(}-\frac{\beta^{3}}{4m}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}$
$\hskip 99.58464pt-\frac{\beta^{4}}{8m^{2}}\big{(}\mathbf{p}\cdot\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{r}})^{2}V(\mathbf{r})\bigg{)}.$ (2.10)

### v3

@@ Abstract
The one-body density matrix (ODM) for a zero temperature non-interacting Fermi gas can be approximately obtained in the semiclassical regime through different $\hbar$-expansion techniques. One would expect that each method of approximating the ODM should yield equivalent density matrices which are both Hermitian and idempotent to any order in $\hbar$. However, the Kirzhnits and Wigner-Kirkwood methods do not yield these properties, while the Grammaticos-Voros method does. Here we show explicitly, for arbitrary $d$-dimensions through an appropriate change into symmetric coordinates, that each method is indeed identical, Hermitian, and idempotent. This change of variables resolves the inconsistencies between the various methods, showing that the non-Hermitian and non-idempotent behaviour of the Kirzhnits and Wigner-Kirkwood methods is an artifact of performing a non-symmetric truncation to the semiclassical $\hbar$-expansions. Our work also provides the first explicit derivation of the $d$-dimensional Grammaticos-Voros ODM, originally proposed by Redjati et al (2019 J. Phys. Chem. Solids 134 313–8) based on their $d=1,2,3,4$ expressions.
Keywords: Density-functional theory, semiclassical, $\hbar$-expansion, Hermitian, Fermi gas
@@ 1 Introduction
Density functional theory (DFT) is a powerful tool used to solve quantum many-body problems. It is applied in various fields of physics at the quantum level such as the calculation of electronic structure and band structures in solid-state physics, and it can also be applied to calculate binding energies of molecules in chemistry (see for instance: [1, 2, 3, 4, 5, 6] and references therein). In DFT, the local particle density $\rho(\mathbf{r})$ constitutes the basic variable to describe the properties of a system of $N$ interacting particles. For a practical use of DFT, Kohn and Sham (KS) proposed to replace the real interacting system with a system of $N$ non-interacting fermions [7, 8]. The ground state of the latter auxiliary system is characterized at every point in space by the same particle density as that of the real system. The KS version of DFT is a one particle theory where each fermion moves in an effective one-body potential $v_{\mathrm{eff}}(\mathbf{r})$. Let $\rho(\mathbf{r})$ be the ground state particle density; in (KS) DFT the non-interacting kinetic energy functional $T_{s}[\rho]=\int\tau(\mathbf{r})d^{d}r$ constitutes an important part of the total-energy, where the superscript $d$ within the integral represents the number of dimensions of the system. Here $\tau(\mathbf{r})$ is the kinetic energy density and is valid for a system at zero absolute temperature [1, 3] in terms of the (KS) single particle states $\{\phi_{j}\}$ with corresponding eigenvalues $\{\varepsilon_{j}\}$, and is given as
$\tau(\mathbf{r})=\frac{\hbar^{2}}{2m}\sum_{\varepsilon_{j}\leq\mu}|\nabla\phi_{j}(\mathbf{r})|^{2},$ (1.1)
where $\mu$ is the Fermi energy. It is possible to express the above positive defined kinetic energy density in terms of the one-body density matrix (ODM), $\rho(\mathbf{r},\mathbf{r}^{\prime})$, given in terms of (KS) single particle states by, $\rho(\mathbf{r},\mathbf{r}^{\prime})=\sum_{\varepsilon_{j}\leq\mu}\phi_{j}(\mathbf{r})\phi_{j}^{\ast}(\mathbf{r}^{\prime})$, so that
$\tau(\mathbf{r})=\frac{\hbar^{2}}{2m}\big{(}\nabla_{\mathbf{r}}\cdot\nabla_{\mathbf{r}^{\prime}}\big{)}\rho(\mathbf{r},\mathbf{r}^{\prime})|_{\mathbf{r}^{\prime}=\mathbf{r}}.$ (1.2)
Notice that the ODM may be seen as the matrix element in coordinate representation of the density operator, $\hat{\rho}=\sum_{\varepsilon_{j}\leq\mu}|\phi_{j}\rangle\langle\phi_{j}|$. The diagonal element $\rho(\mathbf{r},\mathbf{r}^{\prime}=\mathbf{r})\equiv\rho(\mathbf{r})$ is the local density and normalized to the total particle number $N$ of fermions, $\int\rho(\mathbf{r})d^{d}r=N$.
The local density approximation (LDA) is known as the starting point of many approximations of $\rho(\mathbf{r},\mathbf{r}^{\prime})$, and yields the well-known Thomas-Fermi kinetic energy functional [9, 10]. Although this approximation is a good starting point, it often fails to capture the complete physics of a given non-interacting Fermi gas since it neglects the effects of inhomogeneities. For non-uniform systems with finite size, the exact analytical expressions of ODMs and the resulting exact density functionals are in general not known, but density gradient corrections to the LDA exist and can be computed through various techniques  [10, 11, 12, 13, 14]. The exact quantum ODM defined above is Hermitian, $\rho(\mathbf{r},\mathbf{r}^{\prime})=\rho^{\ast}(\mathbf{r}^{\prime},\mathbf{r})$, and idempotent, $\hat{\rho}^{2}=\hat{\rho}$. These properties are satisfied if all terms in the $\hbar$ or gradient expansion of the ODM are accounted for whenever the series converges. A truncation at a given order may lead to a loss of the Hermiticity and idempotency, and clearly this symmetry violation has consequences for quantities that are derived from it [15].
@@
A particular subset of gradient correction techniques includes different semiclassical $\hbar$ expansions, with three popular methods being the Kirzhnitz method (KODM) [16, 17, 18, 19], the Wigner-Kirkwood method (WKODM) [20], and the Grammaticos-Voros method (GVODM) [21]. In reference [22] the KODM was derived up to second order in $\hbar$ for arbitrary $d$ spatial dimensions. Furthermore, in reference [10] the $d$-dimensional WKODM method is discussed although the ODM is not explicitly derived, and in references [23, 24, 25] the GVODM is derived to second order in $\hbar$ for $d=1,2,3,4$-dimensions. Although in each case the ODMs were all derived for a zero temperature non-interacting Fermi gas in the presence of a position-dependent external potential, they can still be used to account for the effects of finite temperature and internal interactions, see references [25, 26].
@@
In reference [22] it was found that when truncated to second order in $\hbar$, the $d$-dimensional KODM and WKODM methods are found to be manifestly non-Hermitian and non-idempotent. In contrast, the GVODM preserves Hermiticity and idempotency regardless of whichever order in $\hbar$ the semiclassical expansion is truncated at. These issues are commented on in references [23, 24, 25]. The various ODMs should all be equivalent since they must yield the same physical properties of the Fermi gas, e.g., the total energy. Therefore it is unsatisfying that the GVODM is manifestly Hermitian and idempotent while the KODM and WKODM are not. Additionally, a $d$-dimensional expression for the GVODM has not yet been explicitly derived. Although reference [23] proposed such an expression based on their $d=1,2,3,4$ results, a rigorous derivation was not performed.
@@
The structure of this paper is as follows: In Section 2, we show that because the WKODM and KODM are formulated in terms of inherently non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$ from the beginning of their derivations, they do not display the same symmetry as the GVODM which is formulated in terms of center of mass $\mathbf{R}$ and relative $\mathbf{s}$ coordinates.
@@
We explicitly show in arbitrary $d$-dimensions that, upon the proper symmetrization procedure (change of coordinates) from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$, the WKODM and KODM are identical to the GVODM up to second order in $\hbar$. Our starting point relies on the ability to express each term of the WKODM and KODM as a specific inverse Laplace transformation before the coordinate transformation is undertaken. We show in A that in terms of non-symmetric coordinates, the WKODM is identical to the KODM via these inverse Laplace transformations [10, 27], allowing us to make the necessary change of coordinates. For completeness, in C we also provide an alternate yet equivalent derivation in $d$-dimensions of the equality between the WKODM (and therefore also KODM) and the GVODM, by making the change of coordinates to $(\mathbf{R},\mathbf{s})$ from the very beginning before any inverse Laplace transformations are evaluated.
The symmetrization procedure for both the KODM and WKODM demonstrates that regardless of whichever method is used to derive the ODM of a zero temperature non-interacting Fermi gas, the same Hermitian and idempotent expression is obtained as one would expect. Our results can be shown to hold when higher order terms in the ODMs are accounted for, but, for simplicity, we restricted our analysis to second order in $\hbar$. This therefore resolves the Hermiticity and idempotency inconsistency between the various methods, showing that problems only arise when non-symmetric truncations of semiclassical expansions are taken. Additionally, this explicitly shows that the predicted $d$-dimensional expression for the GVODM proposed in reference [23] is in fact correct, as we exactly obtain their result through our symmetrization procedure for both the KODM and WKODM methods. To our knowledge, these issues have been overlooked in the semiclassical expansion methods utilized in DFT.
2 Hermitian form of the WKODM and KODM in $d$-dimensions
The $d$-dimensional KODM for a non-interacting Fermi gas in non-symmetric coordinates $(\mathbf{r},\mathbf{r}^{\prime})$ to second order is given by [22]
$\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})=A_{0}^{{K}}+A_{1}^{{K}}+A_{2}^{{K}};$ (2.1)
$A_{0}^{{K}}$ $=\frac{gk_{F}^{d}}{(2\pi z)^{d/2}}J_{\frac{d}{2}}(z),$ (2.2)
$A_{1}^{{K}}$ $=-\frac{gk_{F}^{d-3}z^{2}}{4(2\pi z)^{d/2}}J_{\frac{d}{2}-1}(z)\bigg{(}\nabla_{\mathbf{r}}k_{F}^{2}\cdot\frac{\mathbf{s}}{s}\bigg{)},$ (2.3)
$A_{2}^{{K}}$ $=\frac{gk_{F}^{d-4}z^{2}}{24(2\pi z)^{d/2}}J_{\frac{d}{2}-2}(z)\big{(}\nabla_{\mathbf{r}}^{2}k_{F}^{2}\big{)}+\frac{gk_{F}^{d-6}z^{3}}{96(2\pi z)^{d/2}}J_{\frac{d}{2}-3}(z)\big{(}\nabla_{\mathbf{r}}k_{F}^{2}\big{)}^{2}$
@@ 2 Hermitian form of the WKODM and KODM in $d$-dimensions
The factor $g$ is the spin degeneracy, the local Fermi wavenumber is defined by $k_{F}=k_{F}(\mathbf{r})=\frac{\sqrt{2m}}{\hbar}\big{(}\mu-V(\mathbf{r})\big{)}^{1/2}$ where we have denoted $z=k_{F}s$ with $s=|s|=\sqrt{\mathbf{s}\cdot\mathbf{s}}$, and $J_{\nu}(z)$ is the cylindrical Bessel function of order $\nu$ [28]. For convenience, we have suppressed the $\mathbf{r}$ coordinate dependence of both $k_{F}$ and the $A_{0,1,2}^{{K}}$ terms.
@@
Equation 2.1 with equations 2.2, 2.3 and 2.4 was obtained by simplifying equation (17) from reference [22] using the Bessel function identities described in B, namely equations 2.1 and 2.2. Furthermore, we have collected the various terms of the KODM based on their order of $\hbar$ (or equivalently, their order of gradient of the potential). In equation 2.2, $A_{0}^{{K}}$ is the classical zeroth order term of the semiclassical expansion (order $\hbar^{0}$), also known as the Thomas-Fermi term. It is the only term which does not vanish for uniform systems. To account for effects due to inhomogeneities, one must go beyond the LDA by including density gradient corrections through the terms $A_{1}^{K}$ and $A_{2}^{K}$ which are respectively of order $\hbar$ and 2superscriptPlanck-constant-over-$\hbar^{2}$.
@@
Equation 2.1 is manifestly non-Hermitian and non-idempotent as discussed in references [23, 24, 25]. Since equation 2.1 contains no complex variables and therefore is equal to its own complex conjugate, the non-Hermiticity can be seen by swapping $\mathbf{r}$ and $\mathbf{r}^{\prime}$, equivalent to the replacement $\mathbf{s}\rightarrow-\mathbf{s}$, and observing that the resulting expression is not equal to $\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})$. That is, $\rho_{K}(\mathbf{r},\mathbf{r}^{\prime})\neq\rho_{K}^{\ast}(\mathbf{r}^{\prime},\mathbf{r})$. However, by changing to center of mass $\mathbf{R}$ and relative coordinates $\mathbf{s}$ via equation 1.3, we show that equation 2.1 can be recasted into a manifestly Hermitian and idempotent form identical to the GVODM. The transformation from $(\mathbf{r},\mathbf{r}^{\prime})\rightarrow(\mathbf{R},\mathbf{s})$ variables is non-trivial as we will show. Consequently, to simplify the process, we will individually transform each $A_{0,1,2}^{{K}}$ term in equation 2.1, and sum the results in the end via equation 2.1.
Although in reference [22] equations 2.1, 2.2, 2.3 and 2.4 are derived in $d$-dimensions in terms of non-symmetric coordinates using the commutator method of Kirzhnitz [16], they can equivalently also be derived in these coordinates via the WKODM method [10, 20, 27]. We show this to be true in A so that equations 2.2, 2.3 and 2.4 may alternatively each be expressed as an appropriate inverse Laplace transform of the so-called Bloch density matrix or Bloch propagator. The latter is of great interest in the study of properties of non-interacting fermions in a one-body Hamiltonian $H$. The Bloch density matrix is defined as $C(\mathbf{r},\mathbf{r}^{\prime};\beta)=\langle\mathbf{r}|\textrm{e}^{\beta H}|\mathbf{r}^{\prime}\rangle$, where $\beta$ is a variable which, in general, is taken to be complex, and not the inverse of the temperature. In terms of single particle wavefunctions, the Bloch density matrix reads $C(\mathbf{r},\mathbf{r}^{\prime};\beta)=\sum_{j}\phi_{j}(\mathbf{r})\phi_{j}^{\ast}(\mathbf{r}^{\prime})\textrm{e}^{\beta\varepsilon_{j}}$. By expressing the ODM in terms of the unit step function $\theta(x)$ as $\rho(\mathbf{r},\mathbf{r}^{\prime})=\sum_{j}\phi_{j}(\mathbf{r})\phi_{j}^{\ast}(\mathbf{r}^{\prime})\theta(\mu-\varepsilon_{j})$, where the sum is extended over the whole spectrum, this latter ODM, for a given Fermi energy $\mu$, can be written as an inverse Laplace transform of $C(\mathbf{r},\mathbf{r}^{\prime};\beta)/\beta$ [10]
[…]
Now returning to the semiclassical expansion considered in this work, an expansion in terms of powers of $\hbar$ is known for $C(\mathbf{r},\mathbf{r}^{\prime};\beta)$, and is called the Wigner-Kirkwood expansion [10]. Up to second order we write it as $C^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)\approx C^{WK}_{0}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C^{WK}_{1}(\mathbf{r},\mathbf{r}^{\prime};\beta)+C^{WK}_{2}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ where the first term is the zeroth order term (Thomas-Fermi), the second term is of order $\hbar$, and the third term is of order 2superscriptPlanck-constant-over-$\hbar^{2}$. Substituting the latter $\hbar$-expansion of $C(\mathbf{r},\mathbf{r}^{\prime};\beta)$ into equation 2.5 allows one to obtain the ODM in the Wigner-Kirkwood approach in the form $\rho_{WK}(\mathbf{r},\mathbf{r}^{\prime})=A_{0}^{WK}+A_{1}^{WK}+A_{2}^{WK}$ with
$A_{0}^{{WK}}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{0}^{{WK}}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.6)
$A_{1}^{{WK}}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{1}^{{WK}}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}},$ (2.7)
$A_{2}^{{WK}}=\mathcal{L}^{-1}_{\mu}\bigg{\{}\frac{C_{2}^{{WK}}(\mathbf{r},\mathbf{r}^{\prime};\beta)}{\beta}\bigg{\}}.$ (2.8)
In the above the $C_{0,1,2}^{WK}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ are respectively given by [10]
$C_{0}^{{WK}}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}},$ (2.9)
@@
$C_{1}^{{WK}}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=-\frac{ig\hbar\beta^{2}}{2m(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\Big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{p}\Big{)},$ (2.10)
@@
$C_{2}^{{WK}}(\mathbf{r},\mathbf{r}^{\prime};\beta)$ $=\frac{g\hbar^{2}}{(2\pi\hbar)^{d}}\int d^{d}p\;\textrm{e}^{-\beta H_{cl}(\mathbf{r},\mathbf{p})+\frac{i}{\hbar}\mathbf{p}\cdot\mathbf{s}}\bigg{(}-\frac{\beta^{{2}}}{4m}\nabla_{\mathbf{r}}^{2}V(\mathbf{r})+\frac{\beta^{3}}{6m}\big{(}\nabla_{\mathbf{r}}V(\mathbf{r})\big{)}^{2}$
$\hskip 99.58464pt-\frac{\beta^{4}}{8m^{2}}\big{(}{\nabla_{\mathbf{r}}V(\mathbf{r})\cdot\mathbf{p}}\big{)}^{2}+\frac{\beta^{3}}{6m^{2}}(\mathbf{p}\cdot\nabla_{\mathbf{r}})^{2}V(\mathbf{r})\bigg{)}.$ (2.11)